OpenStudy (anonymous):

Solve 2x2 - 8x = -7.

OpenStudy (anonymous):

x=11/8 if you're solving the equation for x

OpenStudy (anonymous):

its solving quadratic equations

OpenStudy (dmb1215):

is it 2 times 2

OpenStudy (anonymous):

I dont know how to do it. My teacher isnt really around to explain

OpenStudy (dmb1215):

what is it algerbra geometry what

OpenStudy (anonymous):

algebra 2

OpenStudy (dmb1215):

ok be back in a sec

OpenStudy (anonymous):

okay thanks

OpenStudy (dmb1215):

even using quadratic formulas it is 11/8

OpenStudy (anonymous):

it shows it differently not as fractions.

OpenStudy (dmb1215):

whats 11/8

OpenStudy (anonymous):

like -2+ square root of 2

OpenStudy (anonymous):

\[-2 \pm 2\]

OpenStudy (dmb1215):

ok i got yah

OpenStudy (anonymous):

\[-2\pm2\sqrt{2}\]

OpenStudy (anonymous):

its multiple choice..

OpenStudy (dmb1215):

hit me with them

OpenStudy (anonymous):

okay one second

OpenStudy (anonymous):

\[-2\pm \sqrt{2}\]

OpenStudy (anonymous):

\[-2\pm2\sqrt{2}\]

OpenStudy (anonymous):

\[2\pm \sqrt{2} \over 2 \]

OpenStudy (anonymous):

\[2\pm \frac{ \sqrt{2} }{ 2}\]

OpenStudy (anonymous):

Those are all the choice a b c and d

OpenStudy (dmb1215):

um a and b are the same

OpenStudy (dmb1215):

@CGGURUMANJUNATH come help please

OpenStudy (cggurumanjunath):